OpenStudy (nbarrera):

Help+I'll Medal! Express the area of this triangle as a monomial. (info below)

OpenStudy (nbarrera):

A.\[15x ^{2}\] B.\[56x ^{2}\] C.1x D.\[28x ^{2}\]

OpenStudy (anonymous):

not 100% sure but I think its 28

OpenStudy (nbarrera):

or 56 cuz 7*8=56....i guess?

OpenStudy (babyx3boo):

Formula of the area of a triangle is A=1/2bh meaning A=1/2(7)(8) A=28 because 8x7 equals 56 and half of 56 is 28.

OpenStudy (anonymous):

you do 7*8*.5 for triangles

OpenStudy (anonymous):

the awserw is d

OpenStudy (nbarrera):

ooh ok thx